How do you write y=-(x+1)^2-2 in standard and graph it?
Answer by midwood_trail(310)

You were given the parabola in vertex form.
You need to express it in standard form.
The standard form is y = ax^2 + bx + c
Your equation: y = -(x+1)^2 - 2
y = -(x+1)(x+1) - 2
y = -(x^2 + 2x + 1) - 2
y = -x^2 - 2x - 1 - 2
y = -x^2 - 2x - 3
=======================
Since the x^2 term has a negative infront, this means the parabola will be graphed downward (like an up side down letter U) and will have maximum value at the vertex.
To graph, select 3 negative and 3 positive values for x, simplify to find y values that together will make 6 points in the form (x,y). When you have the 6 points, plot them on the xy-plane and then connect each point with a smooth downward curve.
Understand?